5. An electron has a kinetic energy of 6.4 X 10-19 J. What is this energy in electron volts?

 

6. Two metallic plates are 0.04 m apart and the potential difference between them is 50 V. Between the two equipotential surfaces, there is a potential difference of 2.5 V. Find the spacing in this part of the field between two metal plates.


1
Expert's answer
2022-03-27T18:20:57-0400

5.

"6.4*10^{-19}\:\rm J=4\:\rm eV"

6. The electric field between plates

"E=\frac{V}{d}=\frac{50\:\rm V}{0.04\:\rm m}=1250\:\rm V/m"

The spacing between two surfaces

"d'=\frac{V'}{E}=\frac{2.5\:\rm V}{1250\:\rm V/m}=0.002\:\rm m"